OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3)

OpenStudy (chosenmatt):

the first one I think

OpenStudy (anonymous):

thanks :)

OpenStudy (chosenmatt):

no problem

OpenStudy (anonymous):

\(\Huge m=\frac{y_2-y_1}{x_2-x_1}=\frac{1-2}{2-(-3)}\) this is the \(\LARGE \sf SLOPE\) looking at the equation \((y-y_1)=m(x-x_1)\).. i DON'T think the slope (m) is equal to -5.